find the value of iota raise to the power iota

Grade:11

4 Answers

Aman Bansal
12 years ago

Dear Prabhjyot,

cos@ + i.sin@ = e^(i.@)
if we take @ = pi/2:
i = e^(i.pi*/2)
Raise both sides to exponent i:
i^i = e^ (i.pi/2 . i)
i^i = e^(-pi/2)

Best Of luck

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

Priyansh Jain
7 years ago
Answer of i^i is not e^i(-π/2) because its argument as given in my book is 0. But the answer e^i(-π/2) is a contradiction.
Umar
4 years ago
The domain of exponential function is all real numbers.....and real numbers are the subset of complex numbers which means we cannot give ì(iota) as power.......iota is an imaginary quantity so how can we prove of something like imaginary raise to power imaginary...
Vikas TU
3 years ago
Dear student 
i=−1−−−√
i2=(−1−−−√)2=−1
i3=i×i2=i×−1=−i
i4=i2×i2=−1×−1=1
i5=i×i4=i×1=i
i6=i×i5=i×i=i2=−1
i7=i×i6=i×−1=−i
i8=(i2)4=(−1)4=1
i9=i×i8=i×1=i
i10=i×i9=i×i=i2=−1